# When UAV Meets Federated Learning: Latency Minimization via Joint Trajectory Design and Resource Allocation

### v1

@@ Abstract
Federated learning (FL) has emerged as a pivotal solution for training machine learning models over wireless networks, particularly for Internet of Things (IoT) devices with limited computation resources. Despite its benefits, the efficiency of FL is often restricted by the communication quality between IoT devices and the central server. To address this issue, we introduce an innovative approach by deploying an unmanned aerial vehicle (UAV) as a mobile FL server to enhance the training process of FL. By leveraging the UAV’s maneuverability, we establish robust line-of-sight connections with IoT devices, significantly improving communication capacity. To improve the overall training efficiency, we formulate a latency minimization problem by jointly optimizing the bandwidth allocation, computing frequencies, transmit power for both the UAV and IoT devices, and the UAV’s trajectory. Then, an efficient alternating optimization algorithm is developed to solve it efficiently. Furthermore, we analyze the convergence and computational complexity of the proposed algorithm. Finally, numerical results demonstrate that our proposed scheme not only outperforms existing benchmark schemes in terms of latency but also achieves training efficiency that closely approximate the ideal scenario.
@@ I Introduction
Unmanned aerial vehicles (UAVs), also known as drones, have been recently considered as a key technology for the fifth-generation-and-beyond and sixth-generation (B5G/6G) wireless networks and the low-altitude economy [1, 2, 3]. Due to the maneuverability, flexibility, and operability of UAVs, it is believed that the integration of UAVs in B5G/6G wireless networks will provide efficient communication and computation services to the users located in the remote and facility-limited areas [4]. Hence, UAVs have been considered as service providers for various applications, such as military purposes, monitoring stations, delivery of cargoes and medical supplies, and alter-disaster rescues [5].
The deployment of machine learning (ML) algorithms in wireless networks has attracted widespread attention in recent years [6]. Internet of Things (IoT) devices can solve classical wireless problems utilizing advanced ML algorithms, such as communication resource allocation, task offloading, and antenna beamforming design [7, 8]. However, traditional ML algorithms request all devices to upload their local datasets to a remote cloud server for centralized training, which results in high communication overhead and severe privacy breach. To address these issues, federated learning (FL) has emerged as an efficient solution, enabling multiple distributed IoT devices to collaboratively train a shared ML model under the coordination of a central server [9]. In an FL system, each device exchanges model parameters with the central server while retaining raw training data locally, thereby preserving data privacy.
Despite the great potential of FL, its implementation in wireless networks still face a key challenge of unstable transmission quality [10]. Specifically, traditional FL systems take a base station (BS) as the central server to aggregate the model parameters from IoT devices. However, the communication links between IoT devices and the BS may experience blockages, making it challenging for IoT devices to complete local model transmission within a tolerant time threshold. Fortunately, due to the flexible maneuverability and the capability to establish short-distance line-of-sight (LoS) communication links, UAVs can serve as the central server to enable model aggregation and downloading in FL. Nevertheless, both UAVs and IoT devices are typically energy-constrained, while the whole flight and training process require a continuous consumption of energy. Therefore, it is important to optimize the UAV flight trajectory and communication resource allocation to meet the demands of FL training.
@@ II-A UAV-Enabled Wireless Networks
UAV-enabled wireless systems have high efficiency, reliable robustness, and flexible motility [11, 12, 13]. Specifically, Zeng et al. [11] studied a UAV energy consumption minimization problem to improve the communication throughput of mobile users. Xie et al. [12] investigated a minimum throughput maximization problem for all ground users in a wireless powered communication network by optimizing the mobility of the UAV and the wireless resource allocation. Cui et al. [13] studied the dynamic resource allocation with multiple UAVs by utilizing a deep reinforcement learning algorithm. Moreover, the integration of UAVs with other key technologies, such as mobile edge computing (MEC) [14, 15], space-air-ground integrated networks [16, 17], movable-antennas [18], intelligent reflecting surfaces / reconfigurable intelligent surface (IRS/RIS) [19, 20], IoT [21, 22], non-orthogonal multiple access systems [23, 24], blockchain [25], integrated sensing and comminations [26, 27], has significantly enhance the performance of communication systems. These motivate us to introduce UAVs into FL systems to improve both the communication efficiency and training accuracy of FL.
@@ II-C UAV-Supported FL
To overcome the communication bottleneck in FL, several works have utilized the UAV as the central server or mobile relay, to provide LoS links and enhance communication quality [35, 36, 37, 38, 39, 40]. Particularly, Zhang et al. [35] studied the image classification task in a UAV-aided FL system, where a high classification accuracy at relatively low communication cost was achieved. Ng et al. [36] utilized UAVs as mobile relays to facilitate the communications between the vehicles and the server. Pham et al. [37] proposed a UAV-assisted FL system, where the communication resource allocation and the UAV placement were jointly optimized to improve the UAV energy efficiency. Fu et al. [38] optimized the total training time of FL by jointly optimizing the device scheduling, UAV trajectory, and time allocation. UAV-enabled FL systems empowered by convert communications were studied in [39, 40], where UAVs were capable of imposing interfere to the eavesdroppers to further enhance the data security. However, the joint optimization of the UAV trajectory and the allocations of bandwidth, computation resources, and transmit power of both the UAV and IoT devices, has not been fully exploited in previous works.
@@ III System Model
As illustrated in Fig. 1, we consider an FL system that consists of a UAV and $M$ IoT devices. The UAV is equipped with a computing processor for model aggregation, serving as a mobile central server. Each IoT device trains a local ML model and uploads its local model parameters to the UAV. The set of IoT devices is denoted as =$\mathcal{M}=\{1,2,\cdots,M\}$. All devices are fixed within a geographical region of $D\times D\ {\rm{km}}^{2}$. Let $\mathcal{S}_{m}=(\mathbf{s}_{m},0)$ denote the location of the $m$-th IoT device, where $\mathbf{s}_{m}[n]=(x_{m}[n],y_{m}[n])$ denotes the two-dimensional (2D) Cartesian coordinates.
We consider a time block that is divided into $N$ time slots. In each time slot, the UAV needs to perform a round of training, including local model aggregation and global model downloading. The set of time slots is defined as $\mathcal{N}=\{1,2,\cdots,N\}$. Throughout the entire time block, the UAV continues to fly. Therefore, at the beginning of each time slot, we select a denotative position for the UAV to denote its current position. Let ${\mathcal{Q}}[n]=(\mathbf{q}[n],H)$ denote the position of the UAV at the $n$-th time slot, where $\mathbf{q}[n]=(x[n],y[n])$ denotes the 2D Cartesian coordinates and $H_{0}$ is the fixed flying height.
@@ III-A FL Model
We assume that each IoT device has the capability to sense data from its environment to update its local dataset. The number of newly sensed data samples follows a Poisson distribution $\mathcal{P}$ with density $\rho_{\mathrm{sensed}}$, i.e., $D_{m}^{\mathrm{sensed}}[n]=\mathcal{P}(\rho_{\mathrm{sensed}})[n]$ data samples sensed in time slot $n$. Meanwhile, the old data of each IoT device may drop out, where the number of dropped old data samples follows a Poisson distribution with density $\rho_{\mathrm{dropped}}$, i.e., $D_{m}^{\mathrm{dropped}}[n]=\mathcal{P}(\rho_{\mathrm{dropped}})[n]$ data samples dropped in time slot $n$. Therefore, the number of training data samples in time slot $n$ is given by
@@ III-B Local Update and Uploading Model
where $\rho_{0}$ denotes the channel power at the reference distance $d_{0}=1\text{m}$, and $\Delta_{m}[n]$ denotes the distance from the $m$-th IoT device to the UAV in time slot $n$. Moreover, we adopt the OFDMA technology to allocate a total of $U$ orthogonal resource blocks (RBs) to IoT devices. These $U$ RBs are represented by the set $\mathcal{U}=\{1,2,\cdots,U\}$. Let $\alpha_{m,u}[n]=\{0,1\}$ denote the allocation indicator of the $u$-th RB to the $m$-th IoT devices in time slot $n$. Specifically, $\alpha_{m,u}[n]=1$ denotes that the $u$-th RB is allocated to the $m$-th IoT devices, otherwise $\alpha_{m,u}[n]=0$. Therefore, the data rate for the local model uploading can be expressed as
@@ III-D Problem Formulation
$E_{m}^{\max}$ and $E_{\mathrm{UAV}}^{\max}$ represent the maximum energy budgets for the $m$-th IoT device and the UAV, respectively. The constraints (24b)-(24e) represent the kinematic limitations of the UAV, while constraints (24f)-(24h) specify the RB allocation restrictions. Furthermore, constraints (24i) and (24j) represent the computing frequency limitations for IoT devices and the UAV, respectively. constraints (24k) and (24l) represent the transmit power limitations for IoT devices and the UAV, respectively. Constraints (24m) and (24n) establish the energy consumption limits for IoT devices and the UAV, respectively.
@@ IV Joint Optimization of Resource Allocation and UAV trajectory
In this section, we first analyze the convergence of the FL system. Based on this, we propose an efficient AO algorithm to solve the problem (24).
@@ IV-A FL Convergence Analysis
Under Assumption 1 and Assumption 2, the lower bound for the local training rounds, and the lower bound for the total communication rounds for ensuring the convergence of FL are presented in the following Theorem 1.
Theorem 1.
Given the local accuracy $\epsilon_{\mathrm{Local}}$, the local learning rate $\upsilon\leq\frac{2}{L}$, the global accuracy $\epsilon_{\mathrm{UAV}}$, and set $0\leq\varrho\leq\frac{\Upsilon}{L}$, the lower bound of local training rounds $\check{R}$, and the lower bound of global training round $\check{N}$ are given by
$\check{R}\triangleq\frac{2}{\upsilon\Upsilon(2-L\upsilon)}\log_{2}\left(\frac{1}{\epsilon_{\mathrm{Local}}}\right),$ (28)
$\check{N}\triangleq\frac{2L^{2}}{\Upsilon^{2}\varrho(1-\epsilon_{\mathrm{Local}})}\ln\frac{1}{\epsilon_{\mathrm{UAV}}}.$ (29)
@@
Please refer to [31]. ∎
@@ IV-B Problem Transformation
Next, we decompose problem (31) into three subproblems: RB allocation, UAV trajectory optimization, and communication and computation resource allocation.
@@ IV-C RB Allocation
Problem (32) is a MINLP problem with binary variables. To handle this issue, we relax the integer constraints (24f) by
@@
Please refer to Appendix A. ∎
According to (35), the value of $\alpha_{m,u}^{*}[n]$ depends on $\mathcal{W}_{m,u}[n]$ for given the Lagrange multiplier $\boldsymbol{\tilde{\varphi}}$, $\boldsymbol{\tilde{\varpi}}$, $\mathbf{\tilde{\xi}}$, $\mathbf{\tilde{\gamma}}$, and $\mathbf{\tilde{\mu}}$ that are associated with the corresponding constraints (24g)-(24n), and (31b). Here, we provide an explanation for each term in $\mathcal{W}_{m,u}[n]$. The first term is defined as the latency and energy expended by the $m$-th IoT device in selecting the $u$-th RB to upload local model parameters to the UAV. The second term is defined as the latency and energy expended by the UAV in downloading to the $m$-th IoT device over the $u$-th RB.
@@
Accordingly, we utilize the decomposition approach [45] to reconstruct the feasible solution, which is tailored to the considered FL system. During the reconstruction, we divide each RB into $\chi$ sub-RBs. Consequently, the total number of sub-RBs is $U\times\chi$. Thus, the number of sub-RBs occupied by the $m$-th IoT device in time slot $n$ is denoted as $\alpha_{m}^{\rm{total}}[n]=\left\lfloor\chi\times\sum_{u\in\mathcal{U}}\alpha_{m,u}[n]\right\rceil$, where $\lfloor x\rceil$ denotes the nearest integer of $x$.

### v2

# Latency Minimization for UAV-Enabled Federated Learning: Trajectory Design and Resource Allocation

@@ Abstract
Federated learning (FL) has become a transformative paradigm for distributed machine learning across wireless networks. However, the performance of FL is often hindered by the unreliable communication links between resource-constrained Internet of Things (IoT) devices and the central server. To overcome this challenge, we propose a novel framework that employs an unmanned aerial vehicle (UAV) as a mobile server to enhance the FL training process. By capitalizing on the UAV’s mobility, we establish strong line-of-sight connections with IoT devices, thereby enhancing communication reliability and capacity. To maximize training efficiency, we formulate a latency minimization problem that jointly optimizes bandwidth allocation, computing frequencies, transmit power for both the UAV and IoT devices, and the UAV’s flight trajectory. Subsequently, we analyze the required rounds of the IoT devices training and the UAV aggregation for FL convergence. Based on the convergence constraint, we transform the problem into three subproblems and develop an efficient alternating optimization algorithm to solve this problem effectively. Additionally, we provide a thorough analysis of the algorithm’s convergence and computational complexity. Extensive numerical results demonstrate that our proposed scheme not only surpasses existing benchmark schemes in reducing latency up to 15.29%, but also achieves training efficiency that nearly matches the ideal scenario.
@@ I Introduction
Unmanned aerial vehicles (UAVs), also known as drones, are pivotal to sixth-generation (6G) networks and the low-altitude economy [1, 2, 3], offering maneuverability, flexible deployment, and operational efficiency for delivering communication-computation services in remote or infrastructure-limited areas [4]. Their applications span military operations, disaster response, medical logistics, and environmental monitoring [5].
While UAVs have demonstrated significant potential in diverse applications, their seamless integration into 6G wireless networks necessitates advanced strategies to address emerging challenges, such as dynamic resource allocation, stringent latency requirements, and data privacy concerns. To address these issues, machine learning (ML) has emerged as an effective technology [6]. Specifically, the deployment of ML algorithms in wireless networks has attracted widespread attention in recent years [7]. Internet of Things (IoT) devices can solve classical wireless problems utilizing advanced ML algorithms, such as communication resource allocation, task offloading, and antenna beamforming design [8, 9]. However, traditional ML algorithms request all devices to upload their local datasets to a remote cloud server for centralized training, which results in high communication overhead and severe privacy breach. multiple distributed Unlike traditional centralized ML requiring raw data uploads, federated learning (FL) empowers distributed IoT devices to collaboratively train shared models via parameter exchanges with a central server, inherently reducing communication overhead and preserving data privacy [10]. These synergistic capabilities motivate our exploration of UAV-enabled FL systems, leveraging aerial agility for adaptive coverage and privacy-preserving ability of FL to tackle the communication-computation dilemmas in 6G networks.
Despite the great potential of FL, its implementation in wireless networks still face a key challenge of unstable transmission quality [11]. Specifically, traditional FL systems take a base station (BS) as the central server to aggregate the model parameters from IoT devices. However, the communication links between IoT devices and the BS may experience blockages, making it challenging for IoT devices to complete local model transmission within a tolerant time threshold. Fortunately, due to the flexible maneuverability and the capability to establish short-distance line-of-sight (LoS) communication links, the UAVs can serve as the central server to enable model aggregation and downloading in FL. Nevertheless, both the UAVs and the IoT devices are typically energy-constrained, while the whole flight and training process require a continuous consumption of energy. Therefore, it is important to optimize the UAV flight trajectory and communication resource allocation to meet the demands of FL training.
@@ II-A UAV-enabled Wireless Networks
UAV-enabled wireless systems have high efficiency, reliable robustness, and flexible motility [12, 13, 14]. Specifically, Zeng et al. [12] studied a UAV energy consumption minimization problem to improve the communication throughput of mobile users. Xie et al. [13] investigated a minimum throughput maximization problem for all ground users in a wireless powered communication network by optimizing the mobility of the UAV and the wireless resource allocation. Cui et al. [14] studied the dynamic resource allocation with multiple UAVs by utilizing a deep reinforcement learning algorithm. Moreover, the integration of the UAVs with other key technologies, such as mobile edge computing (MEC) [15, 16], space-air-ground integrated networks [17, 18], movable-antennas [19], intelligent reflecting surfaces / reconfigurable intelligent surface (IRS/RIS) [20, 21], IoT [22, 23], non-orthogonal multiple access systems [24, 25], blockchain [26], integrated sensing and comminations [27, 28], has significantly enhance the performance of communication systems. These motivate us to introduce the UAVs into FL systems to improve both the communication efficiency and training accuracy of FL.
@@ II-C UAV-supported FL
To overcome the communication bottleneck in FL, several works have utilized the UAV as the central server or mobile relay, to provide LoS links and enhance communication quality [36, 37, 38, 39, 40, 41]. Particularly, Zhang et al. [36] studied the image classification task in a UAV-aided FL system, where a high classification accuracy at relatively low communication cost was achieved. Ng et al. [37] utilized the UAVs as mobile relays to facilitate the communications between the vehicles and the server. Pham et al. [38] proposed a UAV-assisted FL system, where the communication resource allocation and the UAV placement were jointly optimized to improve the UAV energy efficiency. Fu et al. [39] optimized the total training time of FL by jointly optimizing the device scheduling, the UAV trajectory, and time allocation. UAV-enabled FL systems empowered by convert communications were studied in [40, 41], where the UAVs were capable of imposing interfere to the eavesdroppers to further enhance the data security. However, the joint optimization of UAV trajectory design, bandwidth allocation, computation resource allocation, and transmit power control for both UAV and IoT devices remains underexplored. Existing studies often address these issues separately, failing to achieve the communication-computation-energy trade-off inherent in dynamic air-to-ground networks. Emerging UAV-enabled FL systems demand efficient resource coordination frameworks that holistically optimize these key parameters to ensure energy-efficient FL training while adhering to real-world deployment constraints, such as limited UAV onboard energy, heterogeneous IoT device capabilities, and time-varying channel conditions.
@@ III System Model
As illustrated in Fig. 1, we consider an FL system that consists of a UAV and $M$ IoT devices. Each IoT device continuously collects data to update its local dataset, thereby serving certain privacy-preserving applications, such as surveillance video analysis, and face recognition. The UAV is equipped with a computing processor for model aggregation, serving as a mobile central server. Each IoT device trains a local ML model and uploads its local model parameters to the UAV. The set of IoT devices is denoted as =$\mathcal{M}=\{1,2,\cdots,M\}$. All devices are fixed within a geographical region of $D\times D\ {\rm{km}}^{2}$. Let $\mathcal{S}_{m}=(\mathbf{s}_{m},0)$ denote the location of the $m$-th IoT device, where $\mathbf{s}_{m}[n]=(x_{m}[n],y_{m}[n])$ denotes the two-dimensional (2D) Cartesian coordinates.
We consider a time block that is divided into $N$ time slots. In each time slot, the UAV needs to perform a round of training, including local model aggregation and global model downloading. The set of time slots is defined as $\mathcal{N}=\{1,2,\cdots,N\}$. Throughout the entire time block, the UAV continues to fly. Therefore, at the beginning of each time slot, we select a denotative position for the UAV to denote its current position. Let ${\mathcal{Q}}[n]=(\mathbf{q}[n],H)$ denote the position of the UAV at the $n$-th time slot, where $\mathbf{q}[n]=(x[n],y[n])$ denotes the 2D Cartesian coordinates and $H_{0}$ is the fixed flying height. To enhance readability, the primary notations and their corresponding descriptions are given in Table I.
@@
TABLE I: Notations and Descriptions
Notations Description
$m,M,\mathcal{M}$ The index, number, and set of the IoT devices
$n,N,\mathcal{N}$ The index, number, and set of time slots
$u,U,\mathcal{U}$ The index, number, and set of the RBs
$\mathbf{q}[n]$ The 2D position of the UAV at time slot $n$
$\mathbf{s}_{m}$ The 2D location of the $m$-th IoT device
$D_{m}[n]$ The size of dataset at the $m$-th IoT device
$\mathcal{D}_{m}[n]$ The dataset at the $m$-th IoT device
$\phi_{m}$ The computational resources required for $1$ bit data
$\vartheta$ The data size of one data sample
$R_{m}$ The number of local training rounds of the $m$-th IoT device
$d_{m}[n]$ The size of the local model updates of the $m$-th IoT device
$\rho_{\mathrm{sensed}}$ The Poisson density of sensed data
$\rho_{\mathrm{dropped}}$ The Poisson density of dropped data
$\boldsymbol{\theta}_{m}[n]$ The local model at the $m$-th IoT device
$\boldsymbol{\theta}[n]$ The global model at the UAV
$f_{m}[n]$ The computing frequency of the $m$-th IoT device
$f_{\max}$ The maximum computing frequency of IoT devices
$T_{m}^{\mathrm{train}}[n]$ The local training latency of the $m$-th IoT device
$E_{m}^{\mathrm{train}}[n]$ The energy consumption for training of the $m$-th IoT device
$g_{m}[n]$ The channel gain from the $m$-th IoT device to the UAV
$\alpha_{m,u}[n]$ The allocation indicator of the $u$-th RB to the $m$-th IoT device
$R_{m}^{\rm{up}}[n]$ The data rate for model uploading of the $m$-th IoT device
$T_{m}^{\rm{up}}[n]$ The latency for model uploading of the $m$-th IoT device
$E_{m}^{\rm{up}}[n]$ The energy consumption for model uploading of the $m$-th IoT device
$T_{m}^{\mathrm{wait}}[n]$ The waiting latency of the $m$-th IoT device
$f_{\mathrm{UAV}}[n]$ The computing frequency of the UAV
$F_{\max}$ The maximum computing frequency of the UAV
$T_{\mathrm{UAV}}^{\mathrm{agg}}[n]$ The latency for model aggregation
$E_{\mathrm{UAV}}^{\mathrm{agg}}[n]$ The energy consumption for model aggregation
$T_{m}^{\mathrm{down}}[n]$ The latency of global model downloading for the $m$-th IoT device
$R_{m}^{\mathrm{down}}[n]$ The model downloading rate from the UAV to the $m$-th IoT device
$E_{m}^{\mathrm{down}}[n]$ The energy consumption for model downloading to the $m$-th IoT device
$T_{\mathrm{FL}}[n]$ The overall latency of the $n$-th training round
$L_{m}[n]$ The latency of the $m$-th IoT device
@@ III-A FL Model
We assume that each IoT device has the capability to sense data from its environment to update its local dataset [42]. The number of newly sensed data samples follows a Poisson distribution $\mathcal{P}$ with density $\rho_{\mathrm{sensed}}$, i.e., $D_{m}^{\mathrm{sensed}}[n]=\mathcal{P}(\rho_{\mathrm{sensed}})[n]$ data samples sensed in time slot $n$. Meanwhile, the old data of each IoT device may drop out, where the number of dropped old data samples follows a Poisson distribution with density $\rho_{\mathrm{dropped}}$, i.e., $D_{m}^{\mathrm{dropped}}[n]=\mathcal{P}(\rho_{\mathrm{dropped}})[n]$ data samples dropped in time slot $n$. Therefore, the number of training data samples in time slot $n$ is given by
@@ III-B Local Update and Uploading Model
where $\rho_{0}$ denotes the channel power at the reference distance $d_{0}=1\text{m}$, and $\Delta_{m}[n]$ denotes the distance from the $m$-th IoT device to the UAV in time slot $n$. Moreover, we adopt the OFDMA technology to allocate a total of $U$ orthogonal resource blocks (RBs) to IoT devices. These $U$ RBs are represented by the set $\mathcal{U}=\{1,2,\cdots,U\}$. Let $\alpha_{m,u}[n]=\{0,1\}$ denote the allocation indicator of the $u$-th RB to the $m$-th IoT device in time slot $n$. Specifically, $\alpha_{m,u}[n]=1$ denotes that the $u$-th RB is allocated to the $m$-th IoT device, otherwise $\alpha_{m,u}[n]=0$. Therefore, the data rate for the local model uploading can be expressed as
@@ III-D Problem Formulation
$E_{m}^{\max}$ and $E_{\mathrm{UAV}}^{\max}$ represent the maximum energy budgets for the $m$-th IoT device and the UAV, respectively. The constraints (24b)-(24e) represent the kinematic limitations of the UAV, while constraints (24f)-(24h) specify the RB allocation restrictions. Furthermore, constraints (24i) and (24j) represent the computing frequency limitations for IoT devices and the UAV, respectively. constraints (24k) and (24l) represent the transmit power limitations for IoT devices and the UAV, respectively. Constraints (24m) and (24n) establish the energy consumption limits for IoT devices and the UAV, respectively. The FL training and the constraints among the UAV and the IoT devices are shown in Fig. 2.
Figure 2: The FL training process and the constraints among the UAV and the IoT devices.
@@ IV Joint Optimization of Resource Allocation and UAV Trajectory
In this section, we first discuss the convergence of the FL system. Based on this, we propose an efficient AO algorithm to solve the problem (24).
@@ IV-A FL Convergence Analysis
Under Assumption 1 and 2, the lower bound for the local training rounds, and the lower bound for the total communication rounds for ensuring the convergence of FL are presented in the following Proposition 1 and 2, respectively.
Proposition 1.
Given the local accuracy $\epsilon_{\mathrm{Local}}$, the local learning rate $\upsilon\leq\frac{2}{L}$, the lower bound of local training rounds $\check{R}$ can be given by
$\check{R}=\frac{2}{\upsilon\Upsilon(2-L\upsilon)}\ln\left(\frac{1}{\epsilon_{\mathrm{Local}}}\right).$ (28)
@@
Please refer to Appendix A. ∎
Proposition 2.
Given the global accuracy $\epsilon_{\mathrm{UAV}}$, and set $0\leq\varrho\leq\frac{\Upsilon}{L}$, the lower bound of global UAV aggregation rounds $\check{N}$ can be given by
$\check{N}=\frac{2L^{2}}{\Upsilon^{2}\varrho(1-\epsilon_{\mathrm{Local}})}\ln\left(\frac{1}{\epsilon_{\mathrm{UAV}}}\right).$ (29)
Proof.
Please refer to Appendix B. ∎
@@ IV-B Problem Transformation
Next, we decompose problem (31) into three subproblems: the RB allocation, the UAV trajectory optimization, and communication and computation resource allocation.
@@ IV-C RB Allocation
Problem (32) is an MINLP problem with binary variables. To handle this issue, we relax the integer constraints (24f) by
@@
Please refer to Appendix C. ∎
According to (35), the value of $\alpha_{m,u}^{*}[n]$ depends on $\mathcal{W}_{m,u}[n]$ for given Lagrange multipliers $\boldsymbol{\tilde{\varphi}}$, $\boldsymbol{\tilde{\varpi}}$, $\mathbf{\tilde{\xi}}$, $\mathbf{\tilde{\gamma}}$, and $\mathbf{\tilde{\mu}}$ that are associated with the constraints (24g)-(24n), and (31b). Here, we provide an explanation for each term in $\mathcal{W}_{m,u}[n]$. The first term is defined as the latency and energy expended by the $m$-th IoT device in selecting the $u$-th RB to upload local model parameters to the UAV. The second term is defined as the latency and energy expended by the UAV in downloading to the $m$-th IoT device over the $u$-th RB.
@@
Accordingly, we utilize the decomposition approach [47] to reconstruct the feasible solution, which is tailored to the considered FL system. During the reconstruction, we divide each RB into $\chi$ sub-RBs. Consequently, the total number of sub-RBs is $\tilde{U}=U\times\chi$. Thus, the number of sub-RBs occupied by the $m$-th IoT device in time slot $n$ is denoted as $\alpha_{m}^{\rm{total}}[n]=\left\lfloor\chi\times\sum_{u\in\mathcal{U}}\alpha_{m,u}[n]\right\rceil$, where $\lfloor x\rceil$ denotes the nearest integer of $x$. We note that the reconstruction process may introduce a performance gap. However, it decreases with the bandwidth of the sub-RBs decreases. Therefore, the reconstruction process can always achieve a near-optimal solution to (32) [47].